# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

### v1

@@ 2.5 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
3 Proposed Algorithms
Throughout this paper, we make the following assumptions.
Assumption 3.1.
  1. (i)
    $\phi\in\mathcal{G}(C)$ with $\operatorname{cl}{C}=\operatorname{cl}{\operatorname{dom}\phi}$ is strictly convex on $C\equiv\operatorname{int}\operatorname{dom}\phi$.
  2. (ii)
    $f:\mathbb{R}^{n}\to(-\infty,+\infty]$ is a proper and lower semicontinuous function with $\operatorname{dom}\phi\subset\operatorname{dom}f$, which is $\mathcal{C}^{1}$ on $C$.
  3. (iii)
    The pair $(f,\phi)$ is $L$-smad on $P$.
  4. (iv)
    $P\subset\mathbb{R}^{n}$ is a nonempty compact convex set with $P\subset C$.
Assumption 3.1(i)-(iii) are standard for Bregman-type algorithms [Bolte2018-zt, Takahashi2024-wz]. The strict convexity ensures the existence of $\nu$ by Lemma 2.5 and is satisfied by many kernel generating distances (see Example 2.3). Assumption 3.1(iv) is also standard for FW algorithms, and the condition $P\subset C$ is natural ensuring that the Bregman distance $D_{\phi}$ is well-defined on $P$.
In what follows, let $x^{*}\in\operatorname*{argmin}_{x\in P}f(x)$ and let $f^{*}=f(x^{*})$. We recall a key property for analyzing the convergence rates of FW algorithms in convex optimization.
Lemma 3.2 (Primal gap, dual gap, and Frank–Wolfe gap [Pokutta2024-qu, Lemma 4.1]).
Let $x^{*}\in P$ be an optimal solution of (1.1) and let $f^{*}=f(x^{*})$. Suppose that Assumptions 3.1(ii), (iv) and that $f$ is convex. Then for all $x\in P$, it holds that
3.1 Adaptive Bregman Step-Size Strategy
Assume that $(f,\phi)$ is $L$-smad. Substituting $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}$ with $x_{t},v_{t}\in P$ for $x_{+}$ in Lemma 2.8 provides
$f(x_{t})-f(x_{t+1})\geq\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle-L\gamma_{t}^{1+\nu}D_{\phi}(v_{t},x_{t}).$ (3.2)
Using the optimality condition of the right-hand side of (3.2) in terms of $\gamma_{t}$, we find
$\gamma_{t}=\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}}\quad\text{and}\quad f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle^{1+1/\nu}}{(L(1+\nu)D_{\phi}(v_{t},x_{t}))^{1/\nu}}.$
Theoretically, when $\gamma_{t}\in[0,1]$, the Frank–Wolfe step $x_{t+1}=(1-\gamma_{t})x_{t}+\gamma_{t}v_{t}\in P$ is well-defined. We update $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{\max}\right\}$ with some $\gamma_{\max}\in\mathbb{R}$ (usually set $\gamma_{\max}=1$). This step-size strategy provides the progress lemma. Note that the following lemma does not require the convexity of $f$.
Lemma 3.3 (Progress lemma from the Bregman short step-size).
Proof.
Recalling (3.2), we have
$=\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle,$
If $f$ is convex, using Lemma 3.3, we have
$f(x_{t})-f(x_{t+1})\geq\frac{\nu}{1+\nu}\gamma_{t}\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq\frac{\nu}{1+\nu}\gamma_{t}(f(x_{t})-f^{*}),$
which implies that the primal progress is at least $\frac{\nu}{1+\nu}$ of the Frank–Wolfe gap and of the primal gap multiplied by $\gamma_{t}$. If $\phi=\frac{1}{2}\|\cdot\|^{2}$, we have $\nu=1$ and
$\gamma_{t}=\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|v_{t}-x_{t}\|^{2}},$
which is often called the (Euclidean) short step-size. Although the short step-size strategy does not require line searches, it requires knowledge of the value of $L$.
The exact value or the tight upper bound of $L$ is often unknown, however, the algorithm’s performance heavily depends on it; an underestimation of $L$ might lead to non-convergence, and an overestimation of $L$ might lead to slow convergence. Moreover, a worst-case $L$ might be too conversative for regimes where the function is better behaved. Due to all these reasons, Pedregosa et al. [Pedregosa2020-ay] proposed an adaptive step-size strategy for FW algorithms in the case where $\phi=\frac{1}{2}\|\cdot\|^{2}$. Vyguzov and Stonyakin [Vyguzov2024-tp] proposed a variant using Bregman distances. We have concurrently and independently formulated an improved version similar (but different in details) to [Vyguzov2024-tp], which includes updates to $\nu$ and uses different coefficients for $\gamma_{t}$. We present our algorithm in Algorithm 2.
Output : Updated estimates $\tilde{L}^{*}$ and $\tilde{\nu}^{*}$, step-size $\gamma$
[…]
2      Choose $\beta<1$, $\eta\leq 1$, and $\tau>1$
3     $\kappa\leftarrow 1$
4     $M\leftarrow\eta\tilde{L}$
5     loop
6           $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x),x-v\rangle}{M(1+\kappa)D_{\phi}(v,x)}\right)^{\frac{1}{\kappa}},\gamma_{\max}\right\}$
7          if $f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ then
8                $\tilde{L}^{*}\leftarrow M$
9               $\tilde{\nu}^{*}\leftarrow\kappa$
10               return $\tilde{L}^{*}$, $\tilde{\nu}^{*}$, $\gamma$
11                $M\leftarrow\tau M$
12               $\kappa\leftarrow\beta\kappa$
13
Algorithm 2 Adaptive Bregman step-size strategy
This step-size strategy can be used as a drop-in replacement. For example, inserting $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},1)$ between lines 1 and 1 in Algorithm 1, we obtain that Algorithm 2 searches $L$ and $\nu$ until
$f((1-\gamma)x+\gamma v)-f(x)-\gamma\langle\nabla f(x),v-x\rangle\leq M\gamma^{1+\kappa}D_{\phi}(v,x)$ (3.3)
holds for the approximation values $M$ and $\kappa$ of $L$ and $\nu$, respectively.
Remark 3.4 (Well-definedness and termination of Algorithm 2).
By the $L$-smad property of $(f,\phi)$ and Lemma 2.8, we know that (3.3) holds for all $M\geq L$ and $\nu\geq\kappa>0$. Therefore, the loop in line 2 in Algorithm 2 is well-defined and guaranteed to terminate. Thus it suffices to update $\kappa$ in line 2 only if $D_{\phi}((1-\gamma)x+\gamma v,x)>\gamma^{1+\kappa}D_{\phi}(v,x)$ holds. Thus, $\kappa$ cannot become too small.
3.2 The Away-Step Frank–Wolfe Algorithm
In the case where $P$ is a polytope, the classical FW algorithm might zigzag when approaching the optimal face and, in consequence, converge slowly. To overcome this drawback, Wolfe [Wolfe1970-ax] introduced the away-step Frank–Wolfe algorithm and away steps, which allow the algorithm to move away from vertices in a convex combination of $x_{t}$, which can effectively shortcircuit the zigzagging. The convergence properties of this algorithm were unknown for a long time, and it was only quite recently that Lacoste-Julien and Jaggi [Lacoste-Julien2015-gb] established the linear convergence of the away-step Frank-Wolfe algorithm. Inspired by [GueLat1986-bg, Lacoste-Julien2015-gb, Wolfe1970-ax], we propose a variant of the away-step Frank–Wolfe algorithm utilizing Bregman distances as given in Algorithm 3.
For the following discussion, we introduce the following notions. In the same way [Braun2022-ew], we define an active set $\mathcal{S}\subset\operatorname{Vert}P$ and the away vertex as $v_{t}^{\textup{A}}\in\operatorname*{argmax}_{v\in\mathcal{S}}\langle\nabla f(x_{t}),v\rangle$, where $x_{t}$ is a (strict) convex combination of elements in $\mathcal{S}$, i.e., $x_{t}=\sum_{v\in\mathcal{S}}\lambda_{v}x_{t}$ with $\lambda_{v}>0$ for all $v\in\mathcal{S}$ and $\sum_{v\in\mathcal{S}}\lambda_{v}=1$. Algorithm 3 updates $\nu$ until $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v,x)$ holds in lines 3-3. If $\nu$ is known, one only updates $\gamma_{t}$. If $\nu$ and $L$ are unknown, we can use $L_{t},\nu_{t},\gamma_{t}\leftarrow\textnormal{{step\_size}}(f,\phi,x_{t},v_{t},L_{t-1},\gamma_{t,\max})$ with $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{M(1+\kappa)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\kappa}},\gamma_{t,\max}\right\}$ instead of line 2 in Algorithm 2. When $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$, Algorithm 3 corresponds to the classical away-step FW algorithm.
Input: Initial point $x_{0}\in\operatorname*{argmin}_{v\in P}\langle\nabla f(x),v\rangle$ for $x\in P$, $\beta<1$
1 $\mathcal{S}_{0}\leftarrow\{x_{0}\}$, $\lambda_{x_{0},0}\leftarrow 1$
2for $t=0,\ldots$ do
3      $v_{t}^{\textup{FW}}\leftarrow\operatorname*{argmin}_{v\in P}\langle\nabla f(x_{t}),v\rangle$
4     $v_{t}^{\textup{A}}\leftarrow\operatorname*{argmax}_{v\in\mathcal{S}_{t}}\langle\nabla f(x_{t}),v\rangle$
7           else
8                $v_{t}\leftarrow v_{t}^{\textup{A}}$, $d_{t}\leftarrow v_{t}^{\textup{A}}-x_{t}$, $\gamma_{t,\max}\leftarrow\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$
9                $\nu\leftarrow 1$
10               while $D_{\phi}((1-\gamma)x_{t}+\gamma v_{t},x_{t})\leq\gamma^{1+\nu}D_{\phi}(v_{t},x_{t})$ do
11                     $\gamma\leftarrow\min\left\{\left(\frac{\langle\nabla f(x_{t}),d_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},\gamma_{t,\max}\right\}$, $\nu\leftarrow\beta\nu$
12
13                    $\gamma_{t}\leftarrow\gamma$, $x_{t+1}\leftarrow x_{t}-\gamma_{t}d_{t}$
16                         $\lambda_{v_{t}^{\textup{FW}},t+1}\leftarrow\begin{cases}\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\notin\mathcal{S}_{t}\\ (1-\gamma_{t})\lambda_{v_{t}^{\textup{FW}},t}+\gamma_{t}&\text{if}\ v_{t}^{\textup{FW}}\in\mathcal{S}_{t}\end{cases}$
17                         $\mathcal{S}_{t+1}\leftarrow\begin{cases}\mathcal{S}_{t}\cup\{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}<1\\ \{v_{t}^{\textup{FW}}\}&\text{if}\ \gamma_{t}=1\end{cases}$
18                          else
19                               $\lambda_{v,t+1}\leftarrow(1+\gamma_{t})\lambda_{v,t}$ for all $v_{t}\in\mathcal{S}_{t}\setminus\{v_{t}^{\textup{A}}\}$
20                              $\lambda_{v_{t}^{\textup{A}},t+1}\leftarrow(1+\gamma_{t})\lambda_{v_{t}^{\textup{A}},t}-\gamma_{t}$
22
4 Convergence Analysis for Convex Optimization
In this section, we assume that $f$ is convex.
Assumption 4.1.
The objective function $f$ is convex.
4.1 Sublinear Convergence
We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property. It is a similar result to [Vyguzov2024-tp, Theorem 1]. We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu]. The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
Theorem 4.2 (Primal convergence of the Frank–Wolfe algorithm).
Suppose that Assumptions 3.1 and 4.1 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$. Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$. Then, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (4.1)
and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq\left(\frac{2^{1+\nu}LD^{2}}{\epsilon}\right)^{\frac{1}{\nu}}$.
@@ 4.1 Sublinear Convergence
If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (4.1), we have
$f(x_{t})-f^{*}\leq\frac{4LD^{2}}{t+2},$
which is the same as a sublinear convergence rate of the classical FW algorithm.
Remark 4.3.
While the convergence rate (4.1) is the same as [Vyguzov2024-tp], Vyguzov and Stonyakin assume that the triangle scaling property holds for $D_{\phi}$. In contrast, we require significantly weaker assumptions: it is enough to assume that $\phi$ is strictly convex due to Lemma 2.5 in order to establish Theorem 4.2.
4.2 Accelerated Convergence
We will now establish accelerated convergence rates better than $O(1/t)$ up to linear convergence depending on the choice of parameters. First, we establish the accelerated convergence of Algorithm 1 with the Bregman short step-size, i.e.,
$\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\},$
in the case where the optimal solution lies in the relative interior.
Theorem 4.4 (Accelerated convergence for inner optima).
Suppose that Assumptions 3.1 and 4.1 hold. Let $f$ satisfy the Hölder error bound condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{\frac{1}{\nu}},1\right\}$. Then, it holds that
@@ 4.2 Accelerated Convergence
where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition 2.16, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark 2.13), and $M:=(L(1+\nu))^{1/\nu}$.
@@
This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma 2.7 imply
@@
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim. ∎
Because Algorithm 2 is well-defined (see Remark 3.4), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one from Theorem 4.4 which uses Bregman short steps; up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity. Note that Vyguzov and Stonyakin [Vyguzov2024-tp, Theorem 2] also established linear convergence, assuming that $f$ is relatively strongly convex, the optimal solution belongs to int $P$, the angle condition holds, and the triangle scaling property holds. However, these requirements are stronger than ours here, and they do not establish sublinear but accelerated rates for intermediate parameter regimes.
Next, we establish the linear convergence of Algorithm 3. Recall that the pyramidal width is the minimal $\delta>0$ satisfying Lemma 2.18 (see [Braun2022-ew, Lemma 2.26] or [Lacoste-Julien2015-gb, Theorem 3] for an in-depth discussion). We make the following assumption:
Assumption 4.5.
The kernel generating distance $\phi$ is $\sigma$-strongly convex.
Under Assumption 4.5, it holds that $\delta\leq D$, and this will be important in establishing convergence rates of Algorithm 3.
Theorem 4.6 (Accelerated convergence of the away-step FW algorithm).
Suppose that Assumptions 3.14.1, and 4.5 hold. Let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the Hölder error bound condition with $q>1+\nu$ or $(\nu,q)=(1,2)$. The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}/((\mu/q)^{1/q}\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
@@
By using the induced guarantee on the primal gap via Lemma 2.19, we have
$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (4.2)
@@
where the first inequality holds due to Lemma 3.3 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (4.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
@@
In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps). The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem 4.2 and [Braun2022-ew, Remark 2.4]). We use Lemma 2.20 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim. ∎
Remark 4.7 (Compatibility of parameters).
@@
$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (4.3)
which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough. In order to make (4.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$. When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
5 Convergence Analysis for Nonconvex Optimization
In this section, we consider a nonconvex objective function. We establish global sublinear convergence to a stationary point of (1.1), i.e., $x\in P$ such that $\max_{v\in P}\langle\nabla f(x),x-v\rangle=0$. We also obtain local sublinear and local linear convergence to a minimizer of (1.1).
5.1 Global Convergence

### v3

@@ B.3 Scaling Inequalities, Geometric Hölder Error Bounds, and Contractions
Appendix C Proof of Convergence Analysis for Convex Optimization
C.1 Sublinear Convergence
We establish a sublinear convergence rate of the FW algorithm under the smooth adaptable property. It is a similar result to [Vyguzov2025-ha, Theorem 1]. We conducted its proof following [Braun2022-ew, Jaggi2013-oz, Pokutta2024-qu]. The FW algorithm uses the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$ in the following theorem.
Theorem C.1 (Primal convergence of the Frank–Wolfe algorithm).
Suppose that Assumptions 3.1 and 4.1 hold. Let $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$ characterized by the Bregman distance $D_{\phi}$ and let $\nu\in(0,1]$. Consider the iterates of Algorithm 1 with the open loop step-size, i.e., $\gamma_{t}=\frac{2}{2+t}$. Then, it holds that, for all $t\geq 1$,
$f(x_{t})-f^{*}\leq\frac{2^{1+\nu}LD^{2}}{(t+2)^{\nu}},$ (C.1)
and hence for any accuracy $\epsilon>0$ we have $f(x_{t})-f^{*}\leq\epsilon$ for all $t\geq(2^{1+\nu}LD^{2}/\epsilon)^{1/\nu}$.
@@ C.1 Sublinear Convergence
If $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ in (C.1), we have $f(x_{t})-f^{*}\leq 4LD^{2}/(t+2)$, which is the same as a sublinear convergence rate of the classical FW algorithm.
C.2 Proof of Theorem 4.2
Because Algorithm 2 is well-defined (see Remark 3.5), the convergence result of the FW algorithm using the adaptive step-size strategy (Algorithm 2) is essentially the same as the one that uses Bregman short steps (Theorems 4.2 and 5.2); up to small errors arising from the approximation whose precise analysis we skip for the sake of brevity.
Theorem C.2 (Linear convergence FW algorithm with short step-size or adaptive step-size).
Suppose that Assumptions 3.1 and 4.1 hold. Let $\nu>0$ and let $f$ satisfy the HEB condition with $q\geq 1+\nu$ and $\mu>0$ and $D:=\sqrt{\sup_{x,y\in P}D_{\phi}(x,y)}$ be the diameter of $P$. Assume that there exists a minimizer $x^{*}\in\operatorname{int}P$, i.e., there exists an $r>0$ with $B(x^{*},r)\subset P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\left(\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L(1+\nu)D_{\phi}(v_{t},x_{t})}\right)^{1/\nu},1\right\}$ or Algorithm 2. Then, it holds that, for all $t\geq 1$,
@@
where the second inequality holds from $\langle f(x_{t}),x_{t}-v_{t}\rangle\geq r\|\nabla f(x_{t})\|$ in Proposition B.7, and the last inequality holds because $f$ satisfies the $c$-Łojasiewicz inequality with $c=(q/\mu)^{1/q}$ (see Remark B.4), and $M:=(L(1+\nu))^{1/\nu}$.
@@
This inequality and the initial bound $f(x_{1})-f^{*}\leq LD^{2}$ due to Lemma A.4 imply
@@
Using $f(x_{1})-f^{*}\leq LD^{2}$ and Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=\frac{\nu}{1+\nu}\frac{r^{1+1/\nu}}{Mc^{1+1/\nu}D^{2/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$, we have the claim. ∎
C.3 Proof of Theorem 4.4
Theorem C.3 (Linear convergence of the away-step FW algorithm).
Suppose that Assumptions 3.14.1, and 4.3 hold. Let $\nu>0$ and let $P\subset\mathbb{R}^{n}$ be a polytope and $f$ satisfy the HEB condition with $q>1+\nu$ or $(\nu,q)=(1,2)$. The convergence rate of Algorithm 3 is linear: for all $t\geq 1$
$f(x_{t})-f^{*}\leq\begin{cases}\left(1-\frac{\mu}{32L}\frac{\delta^{2}}{D^{2}}\right)^{\lceil(t-1)/2\rceil}LD^{2}&\text{if}\ (\nu,q)=(1,2),\\ \frac{1}{(1+\nu)^{\lceil(t-1)/2\rceil}}LD^{2}&\text{if}\ 1\leq t\leq t_{0},q>1+\nu,\\ \frac{(L(1+\nu)D^{2}c/(\delta/2)^{1+\nu})^{q/(q-1-\nu)}}{\left(1+\frac{1-\nu}{2(1+\nu)}\lceil(t-t_{0})/2\rceil\right)^{\nu q/(q-1-\nu)}}=\mathcal{O}(1/t^{\nu q/(q-1-\nu)})&\text{if}\ t\geq t_{0},q>1+\nu,\end{cases}$
where $c:=(q/\mu)^{1/q}$ $D$ and $\delta$ are the diameter and the pyramidal width of the polytope $P$, respectively, and
@@
By using the induced guarantee on the primal gap via Lemma B.10, we have
$h_{t}=f(x_{t})-f^{*}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{\langle\nabla f(x_{t}),v_{t}^{\textup{A}}-v_{t}^{\textup{FW}}\rangle}{\delta}\right)^{\frac{q}{q-1}}\leq\left(\frac{q}{\mu}\right)^{\frac{1}{q-1}}\left(\frac{2\langle\nabla f(x_{t}),d_{t}\rangle}{\delta}\right)^{\frac{q}{q-1}},$ (C.2)
@@
where the first inequality holds due to Lemma 3.4 with $\gamma_{t,\max}=1$ (Frank–Wolfe steps) and $\gamma_{t,\max}=\frac{\lambda_{v_{t}^{\textup{A}},t}}{1-\lambda_{v_{t}^{\textup{A}},t}}$ (away steps), and the second inequality holds because of (C.2) and $\langle\nabla f(x_{t}),d_{t}\rangle\geq\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle\geq h_{t}$.
@@
In $q>1+\nu$, we have $h_{t}-h_{t+1}\geq\min\left\{\frac{\nu}{1+\nu},\frac{\nu}{1+\nu}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}h_{t}^{\frac{q-1-\nu}{\nu q}}\right\}\cdot h_{t}$ for at least half of the iterations (Frank–Wolfe steps) and $h_{t+1}\leq h_{t}$ for the rest (away steps). The initial bound $f(x_{1})-f^{*}\leq LD^{2}$ holds generally for the Frank–Wolfe algorithm (see the proof of Theorem C.1 and [Braun2022-ew, Remark 2.4]). We use Lemma B.11 with $c_{0}=LD^{2}$, $c_{1}=\frac{\nu}{1+\nu}$, $c_{2}=c_{1}\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}$, and $\theta_{0}=\frac{q-1-\nu}{\nu q}>0$ from $q>1+\nu$ and obtain the claim. ∎
Remark C.4 (Compatibility of parameters).
@@
$\frac{((\mu/q)^{1/q}\delta/2)^{\frac{1+\nu}{\nu}}}{(L(1+\nu)D^{2})^{1/\nu}}=\left(\frac{\mu}{L}\right)^{1/\nu}\left(\frac{\delta}{D}\right)^{\frac{1+\nu}{\nu}}\frac{1}{2^{\frac{1+\nu}{\nu}}(1+\nu)^{2/\nu}D^{(1-\nu)/\nu}},$ (C.3)
which might be greater than 1 because $\frac{1}{D^{(1-\nu)/\nu}}\geq 1$ if $D<1$ and $\nu$ are small enough. In order to make (C.3) smaller than 1, $\nu$ should be 1, i.e., $(\nu,q)=(1,2)$. When $\nu=1$ and $q\neq 2$, it reduces to the $q>1+\nu$ case.
Appendix D Proof of Convergence Analysis for Nonconvex Optimization
D.1 Global Convergence